# Running gravitational constant induced dark energy as a solution to $\sigma_{8}$ tension

### v1

@@ Abstract
Keywords modified gravity  $\cdot$ scale dependent structure growth  $\cdot$ $\sigma_{8}$ tension release
1 Introduction
@@
To make matters worse, the increased precision of current observations and new analysis techniques have brought to light several tensions between values of different parameters among different observables and experiments. For example, some independent cosmological observations in the redshift range $z\lesssim 0.6$ are in disagreement with Planck18/$\Lambda$CDM. Two of such major tensions are one in the local measurement of the Hubble parameter $H_{0}$ (the so-called Hubble tension), and another in the estimates of the amplitudes of the matter power spectrum on the scales of $8h^{-1}$ Mpc (often referred to as $\sigma_{8}$ tension). Both appear to be in conflict with the corresponding values obtained from Planck18/$\Lambda$CDM. Although the tensions might arise as a result of systematic errors, their persistence after several years of data analysis hints towards a possible breakdown of the standard scenario and the eventual requirement of new physics (see e.g. [8] for a recent review).
In the present article, we shall attempt to address the so-called $\sigma_{8}$ tension. The parameter $\sigma_{8}$ describes the variance of the linear matter perturbations on scales of the order of $8h^{-1}$ Mpc and it can be estimated from several observations, most notably, the weak lensing correlation function obtained by the CFHTLenS collaboration [9], the galaxy cluster count [10] and the Redshift Space Distortion (RSD) measurements [11, 12]. At present there exists a $2\sigma$ tension between the constraints of these local measurements on the matter density $\Omega_{m}$ and the amplitude $\sigma_{8}$ of matter fluctuation, and those from the Planck mission [13, 14]. This tension can be reconciled by modifications to the standard scenario and in the literature, there are several suggestions to alleviate the tension such as different clustering of dark matter at smaller and larger scales [15], a hot dark matter component [16], dark energy-dark matter interaction [17, 18, 19] and modifications of general relativity (GR) [20, 21]. However, many of these solutions are analytically complex, leading to the use of Planck18/$\Lambda$CDM background, or $H(a)$ for simplicity, which can induce undesired bias in the cosmological parameters [22]. On the other hand, there exist some proposed models to alleviate the $H_{0}$ tension, for example with an early dark energy component, which have consequences for the $\sigma_{8}$ tension (see e.g. [23, 24]). It would be preferable that any proposal to alleviate one tension would not make the other worse. In this context, examining simpler, less intrusive modifications becomes essential.
@@
The article is organized as follows: In Section 2.1, we outline the Markov-Mukhanov action and derive the background equations. In Section 2.2, we perturb the model to obtain the equation describing the evolution of the density perturbations. In Section 2.3, we introduce the model-independent parametrization of Newton’s constant and in Section 3.1, we discuss how the model can be incorporated into the standard data analysis methodology. Finally, Section 3.2 presents the findings obtained by fitting the parameterization of $G$ with the data. Throughout the article, we work with the units $c=\hbar=1$.
2 Theory
2.1 Markov-Mukhanov Action
@@ 2.2 Density Perturbations
$G_{\mu}^{\nu}=\bar{G}_{\mu}^{\nu}+\delta G_{\mu}^{\nu}\ ,\quad T_{\mu}^{\nu}=\bar{T}_{\mu}^{\nu}+\delta T_{\mu}^{\nu}\ ,\quad G=\bar{G}+\delta G\ ,\quad\Lambda=\bar{\Lambda}+\delta\Lambda\ .$ (10)
@@
$G_{0}^{0}$ $=a^{-2}\Big{(}3\mathcal{H}^{2}+2\nabla^{2}\Phi-6\mathcal{H}(\Phi^{\prime}+\mathcal{H}\Phi)\Big{)}=a^{-2}\Big{(}3\mathcal{H}^{2}+2\nabla^{2}\Phi\Big{)}\ ,$ (15)
$G^{i}_{i}$ $=-3a^{-2}\left[-\left(2\mathcal{H}^{\prime}+\mathcal{H}^{2}\right)+2\left(\Phi^{\prime\prime}+3\mathcal{H}\Phi^{\prime}+\left(2\mathcal{H}^{\prime}+\mathcal{H}^{2}\right)\Phi\right)\right]\ .$ (16)
@@
$\delta^{{}^{\prime\prime}}_{k}+\mathcal{H}\delta^{\prime}_{k}+\left(\mathcal{H}^{\prime}-\mathcal{H}^{2}\right)\delta_{k}=-2\frac{G^{\prime}}{G}\delta^{\prime}_{k}-\left[\left(1-\frac{k^{2}}{3\mathcal{H}^{2}}\right)\mathcal{H}\frac{G^{\prime}}{G}+\frac{G^{{}^{\prime\prime}}}{G}\right]\delta_{k}\ .$ (24)
@@ 2.3 Parameterizing $G(\rho)$
3 Numerical Analysis
3.1 Method

### v2

@@ Abstract
I Introduction
@@
To make matters worse, the increased precision of current observations and new analysis techniques have brought to light several tensions between values of different parameters among different observables and experiments. For example, some independent cosmological observations in the redshift range $z\lesssim 0.6$ are in disagreement with Planck18/$\Lambda$CDM. Two of such major tensions are one in the local measurement of the Hubble parameter $H_{0}$ (the so-called Hubble tension), and another in the estimates of the amplitudes of the matter power spectrum on the scales of $8h^{-1}$ Mpc (often referred to as $\sigma_{8}$ tension). Both appear to be in conflict with the corresponding values obtained from Planck18/$\Lambda$CDM. On the other hand, recently it has also been shown that $\Omega_{m}$ increases with effective redshift in the late-time data [8, 9], which may pose a challenge to the Planck18/$\Lambda$CDM model. Although the tensions might arise as a result of systematic errors, their persistence after several years of data analysis hints towards a possible breakdown of the standard scenario and the eventual requirement of new physics (see e.g. [10] for a recent review).
In the present article, we shall attempt to address the so-called $\sigma_{8}$ tension. The parameter $\sigma_{8}$ describes the variance of the linear matter perturbations on scales of the order of $8h^{-1}$ Mpc and it can be estimated from several observations, most notably, the weak lensing correlation function obtained by the CFHTLenS collaboration [11], the galaxy cluster count [12] and the Redshift Space Distortion (RSD) measurements [13, 14, 15]. At present there exists a $2\sigma$ tension between the constraints of these local measurements on the matter density $\Omega_{m}$ and the amplitude $\sigma_{8}$ of matter fluctuation, and those from the Planck mission [16, 17]. This tension can be reconciled by modifications to the standard scenario and in the literature, there are several suggestions to alleviate the tension such as different clustering of dark matter at smaller and larger scales [18], a hot dark matter component [19], dark energy-dark matter interaction [20, 21, 22, 23, 24, 25] and modifications of general relativity (GR) [26, 27, 28]. However, many of these solutions are analytically complex, leading to the use of Planck18/$\Lambda$CDM background, or $H(a)$ for simplicity, which can induce undesired bias in the cosmological parameters [29]. On the other hand, there exist some proposed models to alleviate the $H_{0}$ tension, for example with an early dark energy component, which have consequences for the $\sigma_{8}$ tension (see e.g. [30, 31, 32, 33]). It would be preferable that any proposal to alleviate one tension would not make the other worse. In this context, examining simpler, less intrusive modifications becomes essential.
@@
The article is organized as follows: In Section II.1, we outline the Markov-Mukhanov action and derive the background equations. In Section II.2, we perturb the model to obtain the equation describing the evolution of the density perturbations. In Section II.3, we introduce the model-independent parametrization of Newton’s constant and in Section III.1, we discuss how the model can be incorporated into the standard data analysis methodology. Finally, Section III.2 presents the findings obtained by fitting the parameterization of $G$ with the data. Throughout the article, we work with the units $c=\hbar=1$.
II Theory
II.1 Markov-Mukhanov Action
@@ II.2 Density Perturbations
$G_{\mu}^{\nu}=\bar{G}_{\mu}^{\nu}+\delta G_{\mu}^{\nu}\ ,\quad T_{\mu}^{\nu}=\bar{T}_{\mu}^{\nu}+\delta T_{\mu}^{\nu}\ ,\quad$
$G=\bar{G}+\delta G\ ,\quad\Lambda=\bar{\Lambda}+\delta\Lambda\ .$ (10)
@@
$G_{0}^{0}$ $=$ $a^{-2}\Big{(}3\mathcal{H}^{2}+2\nabla^{2}\Phi-6\mathcal{H}(\Phi^{\prime}+\mathcal{H}\Phi)\Big{)}=$ (15)
$=$ $a^{-2}\Big{(}3\mathcal{H}^{2}+2\nabla^{2}\Phi\Big{)}\ ,$
$G^{i}_{i}$ $=$ $-3a^{-2}\left[-\left(2\mathcal{H}^{\prime}+\mathcal{H}^{2}\right)+\right.$ (16)
$+\left.2\left(\Phi^{\prime\prime}+3\mathcal{H}\Phi^{\prime}+\left(2\mathcal{H}^{\prime}+\mathcal{H}^{2}\right)\Phi\right)\right]\ .$
@@
$\delta^{{}^{\prime\prime}}_{k}+\mathcal{H}\delta^{\prime}_{k}+\left(\mathcal{H}^{\prime}-\mathcal{H}^{2}\right)\delta_{k}=$
$=-2\frac{G^{\prime}}{G}\delta^{\prime}_{k}-\left[\left(1-\frac{k^{2}}{3\mathcal{H}^{2}}\right)\mathcal{H}\frac{G^{\prime}}{G}+\frac{G^{{}^{\prime\prime}}}{G}\right]\delta_{k}\ .$ (24)
@@ II.3 Parameterizing $G(\rho)$
III Numerical Analysis
III.1 Method